# Epitome: Pioneering an Experimental Platform for AI-Social Science Integration

### v1

@@ Introduction
Layout and Key Functions By constructing $55$ layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com [^1] http://www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
@@
By constructing $55$ layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com222www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
@@
To demonstrate Epitome’s versatility in facilitating complex human-AI social experiments, we replicated the recent study “Experimental Evidence on the Productivity Effects of Generative Artificial Intelligence” [60], published in Science. This study was replicated using the Epitome platform, conducting a repeated experiment with management professionals while maintaining alignment with the original study’s design and supporting materials (Figure 2.A). We recruited 46 participants who self-identified as experienced, college-educated managers (Dataset S1). Eleven participants were excluded from the analysis due to misreporting their occupation or failing to meet the specified time requirements in the exact-time group, resulting in a final sample of 35 participants. Participants were randomly assigned to either the treatment or control group. Each participant completed two incentivized writing tasks relevant to managerial roles, designed to simulate real-world responsibilities such as drafting performance evaluations and composing strategic communications. Between the two tasks, the treatment group was instructed to register for ChatGPT and received guidance on its use, whereas the control group was asked to register for the LaTeX editor Overleaf, ensuring comparable time and effort investment in learning a new tool. To ensure methodological rigor, all experimental procedures—including task instructions, incentive structures, evaluation criteria, and data collection protocols—were meticulously replicated from the original study(see Figs. S1–S7 for the detailed application of the experimental design).
@@
Case study 2: Human Creativity and Generative AI
Figure 5: Visual representation of Case Study 3. (A)Case 3 investigates AI-driven role interaction in moral dilemma scenarios. Seven fictional characters were developed and introduced in the “AI Town” simulation to model AI-to-AI behavior. These character setting were then converted into prompts and deployed in a multi-agent chatroom, where additional AI agents participated in extended social interactions. The participants engaged in courtroom-style simulations, interacting with AI agents to explore moral decision-making and ethical dynamics in a controlled environment. (B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 3). (C)The AI Town multi-agent simulation environment and the multi-agent chatroom broaden research methodologies by enabling comparative studies of human versus AI decision-making in moral dilemmas for Case Study 3, thereby advancing social science theory.
In the second experiment, we replicate the study published in Science, titled “Generative AI enhances individual creativity but reduces the collective diversity of novel content” [61]. The purpose of the replication is to assess how Epitome can support high-fidelity replication of creative tasks and to extend the original findings by testing them in a multilingual setting (Chinese) and evaluating AI’s impact on both individual creativity and collective diversity of outputs.
The original experimental design involved three groups: a control group (Human-only), a group with one AI-generated suggestion, and a group with five AI-generated suggestions. We replicated the task structure, randomization process, and evaluation criteria. The only modification was the introduction of a Chinese language condition to assess the cross-cultural generalizability of the original study’s findings. The multilingual compatibility of Epitome effectively supports the broad recruitment of participants from diverse language backgrounds. Epitome’s ‘My Material’ and ‘My Bots’ modules were used to ensure that the AI interactions were faithfully replicated, and participants were randomly assigned to one of the three experimental conditions (see Figs. S11–S21 for the detailed application of the experimental design)..
@@
Epitome provides valuable insights into the non-linear effects of AI input on creativity. Specifically, while a single AI suggestion improved output quality, five suggestions led to a reduction in diversity, confirming diminishing returns on AI assistance. The platform’s real-time tracking of AI interactions enabled detailed analysis of how AI-generated ideas influence creative work. This feature, along with the platform’s flexibility in adjusting the number and type of AI suggestions, offered a nuanced understanding of AI’s impact on creativity. In the replication study, participants were randomly assigned to one of three groups—no AI assistance, one AI suggestion, or five AI suggestions—and to one of three creative themes (jungle, planet, or ocean). This design, which required the deployment of multiple agents across different scenarios, significantly increased experimental complexity. To manage this, we utilized Epitome’s “My Workflow” module to create a multi-stage dialogue workflow, enabling precise processing of user inputs and generating targeted creative outputs. Epitome’s integrated capabilities not only streamlined the experimental process but also provided deeper insights into how varying levels of AI assistance shape creative outcomes, contributing significantly to our understanding of AI’s role in human creativity.
@@
We utilized two Epitome modules for the experimental design. In the first phase, using Epitome’s AI Town module, we created an AI community based on the Heinz Dilemma framework and developed seven AI characters with distinct backgrounds, motivations, and objectives, based on roles from the original experiment. These AI characters were then placed within the community to make decisions reflecting moral judgments (see SI.6 for detailed role settings). In the second phase, we expanded the experiment into a specific scenario using Epitome’s multi-agent chatroom module, where LLM-based AI agents engaged in interactions with human participants(see SI.7 for detailed role settings). This phase incorporated a simulated trial scenario, inspired by the common law judicial process, in which participants were assigned roles as judges, jurors, and witnesses, each representing different perspectives on the ethical dilemma. Participants engaged in the trial process in the chatroom, debating with AI agents and performing a series of tasks, ultimately rendering a verdict on whether Heinz was guilty or not. Through Epitome’s real-time interactions and comprehensive logging, we collected valuable data that aids in analyzing the different strategies and preferences employed by humans and AI agents in handling complex moral issues(see Figs. S30–S35 for the detailed application of the experimental design).
@@
The AI Town simulation showed that AI agents consistently favored justice, even when Heinz attempted to defy the law in the fourth simulation, with the narrative remaining focused on justice-related themes, as confirmed by the LDA analysis (see Table S15). In the courtroom simulation, AI judges closely aligned with human judges. However, differences emerged in the roles of jurors and witnesses. Female human participants were more inclined towards care ethics, while male participants and AI agents prioritized justice. AI’s emotional responses remained stable and positive in both simulations, with sentiment scores averaging around 7/10, indicating minimal fluctuation and sustained positivity (Figs. 36a-b). Throughout the four iterations, AI consistently favored justice. In the first three iterations, Heinz, persuaded by his wife, chose actions aligned with justice (see Table S14). Even in the fourth iteration, when Heinz resisted, the LDA analysis confirmed AI’s continued preference for justice. In the courtroom, both AI judges and human judges favored a “guilty but lenient” verdict 87.5% and 75% of the time, respectively (Figs. 36c). Female judges showed a stronger preference for leniency, with 100% of female participants choosing this option compared to 50% of male participants. When acting as jurors, AI and human participants both leaned toward justice, though female human participants showed more variability, balancing justice with care (Figs. 36d). These results highlight the contrast in moral reasoning: human participants, especially women, are more influenced by care ethics, while AI remains focused on justice and legal norms. This aligns with previous studies on gendered moral priorities and extends the framework of AI-assisted adjudication proposed by de Sousa et al. . Moreover, the emotional stability of AI’s decisions contrasts with human emotional sensitivity, underscoring the difference between human and AI decision-making.
@@ Discussion
In terms of technical capabilities, Epitome delivers four key advantages: first, its modular architecture and flexible interface definitions allow seamless integration of diverse large language models, extensible simulation environments, and custom experimental scenarios; second, by embedding an intelligent agent workflow engine with dynamic load-balancing, the platform ensures efficient resource allocation and robust performance under high-concurrency LLM inference; third, a personalized database and knowledge-base management system provides full traceability of experimental data and knowledge artifacts, underpinning rigorous social-science data analysis, reproducibility, and validation; and fourth, dual support for low-code configuration and operator “plug-and-play” development drastically lowers the barrier to experiment design and deployment, accelerating iteration cycles and fostering innovation in human–AI interaction studies.
@@
Data, Materials, and Software Availability
Purpose: Provide the URL of the original dataset for easy access and use by other researchers, along with an open access statement.
ACKNOWLEDGMENTS
Please include your acknowledgments here, set in a single paragraph. Please do not include any acknowledgments in the Supporting Information, or anywhere else in the manuscript.

### v2

@@ Introduction
Layout and Key Functions By constructing 555 layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com [^1] http://www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
@@
By constructing 555 layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com222www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
@@
To demonstrate Epitome’s versatility in facilitating complex human-AI social experiments, we replicated the recent study "Experimental Evidence on the Productivity Effects of Generative Artificial Intelligence" [60], published in Science. This study was replicated using the Epitome platform, conducting a repeated experiment with management professionals while maintaining alignment with the original study’s design and supporting materials (Figure 3). We recruited 46 participants who self-identified as experienced, college-educated managers (Dataset S1). Eleven participants were excluded from the analysis due to misreporting their occupation or failing to meet the specified time requirements in the exact-time group, resulting in a final sample of 35 participants. Participants were randomly assigned to either the treatment or control group. Each participant completed two incentivized writing tasks relevant to managerial roles, designed to simulate real-world responsibilities such as drafting performance evaluations and composing strategic communications. Between the two tasks, the treatment group was instructed to register for ChatGPT and received guidance on its use, whereas the control group was asked to register for the LaTeX editor Overleaf, ensuring comparable time and effort investment in learning a new tool. To ensure methodological rigor, all experimental procedures—including task instructions, incentive structures, evaluation criteria, and data collection protocols—were meticulously replicated from the original study(see Figs. S1–S7 for the detailed application of the experimental design).
@@
In the second experiment, we replicate the study published in Science, titled "Generative AI enhances individual creativity but reduces the collective diversity of novel content" [61]. The purpose of the replication is to assess how Epitome can support high-fidelity replication of creative tasks and to extend the original findings by testing them in a multilingual setting (Chinese) and evaluating AI’s impact on both individual creativity and collective diversity of outputs.
The original experimental design involved three groups: a control group (Human-only), a group with one AI-generated suggestion, and a group with five AI-generated suggestions. We replicated the task structure, randomization process, and evaluation criteria. The only modification was the introduction of a Chinese language condition to assess the cross-cultural generalizability of the original study’s findings. The multilingual compatibility of Epitome effectively supports the broad recruitment of participants from diverse language backgrounds(Figure 4). Epitome’s ‘My Material’ and ‘My Bots’ modules were used to ensure that the AI interactions were faithfully replicated, and participants were randomly assigned to one of the three experimental conditions (see Figs. S11–S21 for the detailed application of the experimental design)..
@@
Epitome provides valuable insights into the non-linear effects of AI input on creativity. Specifically, while a single AI suggestion improved output quality, five suggestions led to a reduction in diversity, confirming diminishing returns on AI assistance. The platform’s real-time tracking of AI interactions enabled detailed analysis of how AI-generated ideas influence creative work. This feature, along with the platform’s flexibility in adjusting the number and type of AI suggestions, offered a nuanced understanding of AI’s impact on creativity. In the replication study, participants were randomly assigned to one of three groups—no AI assistance, one AI suggestion, or five AI suggestions—and to one of three creative themes (jungle, planet, or ocean). This design, which required the deployment of multiple agents across different scenarios, significantly increased experimental complexity. To manage this, we utilized Epitome’s "My Workflow" module to create a multi-stage dialogue workflow, enabling precise processing of user inputs and generating targeted creative outputs. Epitome’s integrated capabilities not only streamlined the experimental process but also provided deeper insights into how varying levels of AI assistance shape creative outcomes, contributing significantly to our understanding of AI’s role in human creativity.
@@
Figure 5: Visual representation of Case Study 3. (A)Case 3 investigates AI-driven role interaction in moral dilemma scenarios. Seven fictional characters were developed and introduced in the "AI Town" simulation to model AI-to-AI behavior. These character setting were then converted into prompts and deployed in a multi-agent chatroom, where additional AI agents participated in extended social interactions. The participants engaged in courtroom-style simulations, interacting with AI agents to explore moral decision-making and ethical dynamics in a controlled environment. (B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 3). (C)The AI Town multi-agent simulation environment and the multi-agent chatroom broaden research methodologies by enabling comparative studies of human versus AI decision-making in moral dilemmas for Case Study 3, thereby advancing social science theory.
@@
We utilized two Epitome modules for the experimental design(Figure 5). In the first phase, using Epitome’s AI Town module, we created an AI community based on the Heinz Dilemma framework and developed seven AI characters with distinct backgrounds, motivations, and objectives, based on roles from the original experiment. These AI characters were then placed within the community to make decisions reflecting moral judgments (see SI.6 for detailed role settings). In the second phase, we expanded the experiment into a specific scenario using Epitome’s multi-agent chatroom module, where LLM-based AI agents engaged in interactions with human participants(see SI.7 for detailed role settings). This phase incorporated a simulated trial scenario, inspired by the common law judicial process, in which participants were assigned roles as judges, jurors, and witnesses, each representing different perspectives on the ethical dilemma. Participants engaged in the trial process in the chatroom, debating with AI agents and performing a series of tasks, ultimately rendering a verdict on whether Heinz was guilty or not. Through Epitome’s real-time interactions and comprehensive logging, we collected valuable data that aids in analyzing the different strategies and preferences employed by humans and AI agents in handling complex moral issues(see Figs. S30–S35 for the detailed application of the experimental design).
@@
The AI Town simulation showed that AI agents consistently favored justice, even when Heinz attempted to defy the law in the fourth simulation, with the narrative remaining focused on justice-related themes, as confirmed by the LDA analysis (see Table S15). In the courtroom simulation, AI judges closely aligned with human judges. However, differences emerged in the roles of jurors and witnesses. Female human participants were more inclined towards care ethics, while male participants and AI agents prioritized justice. AI’s emotional responses remained stable and positive in both simulations, with sentiment scores averaging around 7/10, indicating minimal fluctuation and sustained positivity (Figs. 36a-b). Throughout the four iterations, AI consistently favored justice. In the first three iterations, Heinz, persuaded by his wife, chose actions aligned with justice (see Table S14). Even in the fourth iteration, when Heinz resisted, the LDA analysis confirmed AI’s continued preference for justice. In the courtroom, both AI judges and human judges favored a "guilty but lenient" verdict 87.5% and 75% of the time, respectively (Figs. 36c). Female judges showed a stronger preference for leniency, with 100% of female participants choosing this option compared to 50% of male participants. When acting as jurors, AI and human participants both leaned toward justice, though female human participants showed more variability, balancing justice with care (Figs. 36d). These results highlight the contrast in moral reasoning: human participants, especially women, are more influenced by care ethics, while AI remains focused on justice and legal norms. This aligns with previous studies on gendered moral priorities and extends the framework of AI-assisted adjudication proposed by de Sousa et al.. Moreover, the emotional stability of AI’s decisions contrasts with human emotional sensitivity, underscoring the difference between human and AI decision-making.
@@ Discussion
In terms of technical capabilities, Epitome delivers four key advantages: first, its modular architecture and flexible interface definitions allow seamless integration of diverse large language models, extensible simulation environments, and custom experimental scenarios; second, by embedding an intelligent agent workflow engine with dynamic load-balancing, the platform ensures efficient resource allocation and robust performance under high-concurrency LLM inference; third, a personalized database and knowledge-base management system provides full traceability of experimental data and knowledge artifacts, underpinning rigorous social-science data analysis, reproducibility, and validation; and fourth, dual support for low-code configuration and operator "plug-and-play" development drastically lowers the barrier to experiment design and deployment, accelerating iteration cycles and fostering innovation in human–AI interaction studies.